# Strong constraints on a sharp change in $G$ as a solution to the Hubble tension

### v1

@@ Acknowledgements
IB and HD are supported by Royal Society University Research Fellowship grant 211046. NS is supported by the AKTION grant number MPC-2023-07755, the Charles University Grants Agency (GAUK) grant number 94224, and the Deutscher Akademischer Austauschdienst Bonn-Prague exchange program. The authors thank Earl Bellinger and Jeremy Sakstein for helpful discussions. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Additional consequences of the GSM & future tests
The GSM would affect the propagation of gravitational waves (GWs; LIGO Collaboration, 2016). It has been argued that GWs would propagate continuously, even during the period in which $G$ varies rapidly (Paraskevas & Perivolaropoulos, 2023). Nonetheless, it seems inevitable that the GSM would have different effects on distances based on stellar populations and those obtained using GWs as standard sirens. Further progress is possible using GWs with an electromagnetic counterpart, with one such instance having been reported so far (GW170817; see Virgo & LIGO Collaborations, 2017). Its distance of 40 Mpc is right where the transition is proposed to have occurred. In the future, it should be possible to find similar events either side of the transition. This might eventually allow a determination of the maximum mass of neutron stars, which is sensitive to $G$ and would therefore differ either side of the transition (Reyes & Sakstein, 2024).
@@
If the proposed reduction in $G$ occurred on a short enough timescale, any bound orbits would be somewhat destabilised. A sudden drop in the equilibrium circular velocity would cause an object initially on a circular orbit to find itself on a wider elliptical orbit. In the case of the Earth, this situation would arise if the transition occurred over $\ll 1$ yr. Its eccentricity would be ‘pumped’ by a sharp change in $G$. This is disfavoured by the fact that Earth’s orbital eccentricity is only 0.017 and that of Venus only 0.007.
There would also be consequences elsewhere in the Solar System, including on the giant planets. In general, the whole Solar System would be somewhat destabilised. This would increase the likelihood of giant impacts on the Earth. Indeed, this very possibility has been related to the extinction of the dinosaurs and many other species in the Cretaceous–Paleogene extinction (Perivolaropoulos, 2022), which occurred 66 Myr ago (Renne et al., 2013). This corresponds to a distance of only 20 Mpc, but there are Type Ia SNe with Cepheid calibration out to $\approx 40$ Mpc (e.g., Riess et al., 2022a). Even so, recent studies suggest that the GSM can be reconciled with a transition at a lookback time corresponding to only 20 Mpc (Ruchika et al., 2024a, b). It therefore remains possible that the GSM is associated with the Cretaceous–Paleogene extinction event, though it is not presently known whether the asteroid impact largely responsible for it was an isolated incident or part of a more general increase in giant impacts at that time, as would be expected in the GSM.
@@
Finally, returning to the smaller scale of the Earth, a change in $G$ would affect how it maintains hydrostatic equilibrium. The proposed reduction in $G$ would cause it to expand slightly, as overlying layers of rock exert less weight. This may lead to unusual tectonic effects, possibly triggering earthquakes and volcanism. Life on the Earth would have to respond to these effects, and also more directly to the reduced surface gravity. Since complex life was widespread by the time of the proposed transition, it may have had interesting consequences for especially large land animals like dinosaurs -- unless the transition is associated with their demise. This of course depends on the duration of the transition, which we have generally assumed would occur over $\gg 1$ yr. Life would undoubtedly adapt to a more gradual change in $G$ over several kyr or Myr, though there might be interesting evolutionary consequences.

### v2

@@ Acknowledgements
IB and HD are supported by Royal Society University Research Fellowship grant 211046. NS is supported by the AKTION grant number MPC-2023-07755, the Charles University Grants Agency (GAUK) grant number 94224, and the Deutscher Akademischer Austauschdienst Bonn-Prague exchange program. The authors thank Earl Bellinger, Charles Dalang, Hugo Moreira and Jeremy Sakstein for helpful discussions. They also thank the referee for comments which greatly improved this contribution. For the purpose of open access, the authors have applied a Creative Commons Attribution (CC BY) licence to any Author Accepted Manuscript version arising.
@@ Appendix A Additional consequences of the GSM & future tests
The distance ladder can be calibrated with many different types of star, whose luminosities exhibit a variety of dependencies on the value of $G$. Thus, distance indicators calibrated to match each other within 40 Mpc would likely become discrepant beyond the proposed transition. For instance, Cepheid variable pulsations respond differently to main sequence stars (Jain et al., 2013; Sakstein, 2013), whose average luminosity is essentially what the SBF technique measures (Cantiello & Blakeslee, 2023). The TRGB magnitude (Anand et al., 2022, 2024) responds still differently (Sakstein et al., 2019). Geometric distances to megamasers would not be much affected (Pesce et al., 2020); as all but one of the megamasers considered by those authors are beyond the GSM transition, further constraints could be obtained by comparing megamaser distances to those obtained in ways that are sensitive to $G$.
The GSM would also affect the propagation of gravitational waves (GWs; LIGO Collaboration, 2016). It has been argued that GWs would propagate continuously, even during the period in which $G$ varies rapidly (Paraskevas & Perivolaropoulos, 2023). Nonetheless, it seems inevitable that the GSM would have different effects on distances based on stellar populations and those obtained using GWs as standard sirens. Further progress is possible using GWs with an electromagnetic counterpart, with one such instance having been reported so far (GW170817; see Virgo & LIGO Collaborations, 2017). Its distance of 40 Mpc is right where the transition is proposed to have occurred. In the future, it should be possible to find similar events either side of the transition. This might eventually allow a determination of the maximum mass of neutron stars, which is sensitive to $G$ and would therefore differ either side of the transition (Reyes & Sakstein, 2024). The GSM would also lead to a significant change in the gravitational binding energy of any compact object, with neutron stars particularly severely affected (Goldman, 2024).
@@
If the proposed reduction in $G$ occurred on a short enough timescale, any bound orbits would be somewhat destabilised. A sudden drop in the equilibrium circular velocity would cause an object initially on a circular orbit to find itself on a wider elliptical orbit. In the case of the Earth, this situation would arise if the transition occurred over $\ll 1$ yr. Its eccentricity would be ‘pumped’ by a sharp change in $G$. This is disfavoured by the fact that Earth’s orbital eccentricity is only 0.017 and that of Venus is only 0.007.
There would also be consequences elsewhere in the Solar System, including on the giant planets. In general, the whole Solar System would be somewhat destabilised. This would increase the likelihood of giant impacts on the Earth. Indeed, this very possibility has been related to the extinction of the dinosaurs and many other species in the Cretaceous–Paleogene extinction (Perivolaropoulos, 2022), which occurred 66 Myr ago (Renne et al., 2013). This corresponds to a distance of only 20 Mpc, but there are Type Ia SNe with Cepheid calibration out to $\approx 40$ Mpc (e.g., Riess et al., 2022a). Even so, recent studies suggest that the GSM can be reconciled with a transition at a lookback time corresponding to only 20 Mpc (Ruchika et al., 2024a, b). It therefore remains possible that the GSM is associated with the Cretaceous–Paleogene extinction event, though this is disfavoured by geochronometric constraints on the number of days per year a few Myr prior to this very clear geological boundary (see figure 7 of de Winter et al., 2020). It is not presently known whether the asteroid impact largely responsible for it was an isolated incident or part of a more general increase in giant impacts at that time, as would be expected in the GSM.
@@
Finally, returning to the smaller scale of the Earth, a change in $G$ would affect how it maintains hydrostatic equilibrium. The proposed reduction in $G$ would cause it to expand slightly, as overlying layers of rock exert less weight. This may lead to unusual tectonic effects, possibly triggering earthquakes and volcanism. Life on the Earth would have to respond to these effects, and also more directly to the reduced Solar radiation and surface gravity. Since complex life was widespread by the time of the proposed transition, it may have had interesting consequences for especially large land animals like dinosaurs -- unless the transition is associated with their demise. This of course depends on the duration of the transition, which we have generally assumed would occur over $\gg 1$ yr. Life would undoubtedly adapt to a more gradual change in $G$ over several kyr or Myr, though there might be interesting evolutionary consequences.